OpenStudy (anonymous):

hey i need help urgently .... evaluate integral ∫5 4 8x(cube) dx

OpenStudy (anonymous):

54*8\[54*8 (x^4)/4\]

OpenStudy (anonymous):

Do you mean: \[\int\limits_{4}^{5}8x ^{3}\]

OpenStudy (anonymous):

Yeah I'm not sure what your constants are in this problem

OpenStudy (anonymous):

if so, it is: 738

OpenStudy (anonymous):

The real thing to look at here is that it's really a simple power rule problem... Ax^n+1 / n really simple when you look at it like that

OpenStudy (anonymous):

\[\int\limits_{4}^{5}8x ^{3}dx=\left[ (8/4)x ^{4} \right]_{4}^{5}=\left[ (8/4)(5^{4}) \right]-\left[ (8/4)(4^{4}) \right]\]

OpenStudy (anonymous):

you're right Tfraiz, but since there's someone who asked about it, it means they are having trouble solving it :), we are here to help with details :)

OpenStudy (anonymous):

Hey thanks evryone for such a helpful efforts

OpenStudy (anonymous):

so the answer will be 738 right ?

OpenStudy (anonymous):

If its a definite integral with the bounds from 4 to 5 then yeah it will be 738

OpenStudy (anonymous):

ok